10 6 0 4 8 12 16 20

a) calculate the slope across arc AB, assuming that at point A x = 2 and y = 9 and that point B x= 4 and y = 6

b) calculate the slope across arc BC, assuming that at point B x = 4 and y = 6 and that at point C x = 8 and y = 3

C) calculate the slope across arc CD , assuming that at point C x = 8 and y=3 and that at point D x= 16 and y = 0.
Answer #1

a) Slope arc AB = (y2-y1)/(x2-x1)= (6-9)/(4-2) =-1.5

b) Slope arc BC = (y2-y1)/(x2-x1)= (3-6)/(8-4) =-0.75

c) Slope arc CD = (y2-y1)/(x2-x1)= (0-3)/(16-8) =-0.375